A projectile is launched from the north pole with
Answer #1

h = 0.61 R

R+h = 1.61R

g =GM/R^2 = 9.825 m/s^2, R = 6371 km

From conservation of energy

(P.E+K.E)at earth surface = (P.E+K.E) at maximum altitude

-GMm/R +(1/2)mvo^2 = -GMm/(R+h) +0

-gR +(1/2)vo^2 = -GM/1.61R

-gR+(1/2)vo^2 = -gR/1.61

vo^2 = 2gR[1-1/1.61] =2*9.825*6371*1000*0.379

vo= 6887 m/s